### v1

# Ask in Any Modality: A Comprehensive Survey on Multimodal Retrieval-Augmented Generation

Abstract
Large Language Models (LLMs) struggle with hallucinations and outdated knowledge due to their reliance on static training data. Retrieval-Augmented Generation (RAG) mitigates these issues by integrating external dynamic information enhancing factual and updated grounding. Recent advances in multimodal learning have led to the development of Multimodal RAG, incorporating multiple modalities such as text, images, audio, and video to enhance the generated outputs. However, cross-modal alignment and reasoning introduce unique challenges to Multimodal RAG, distinguishing it from traditional unimodal RAG. This survey offers a structured and comprehensive analysis of Multimodal RAG systems, covering datasets, metrics, benchmarks, evaluation, methodologies, and innovations in retrieval, fusion, augmentation, and generation. We precisely review training strategies, robustness enhancements, and loss functions, while also exploring the diverse Multimodal RAG scenarios. Furthermore, we discuss open challenges and future research directions to support advancements in this evolving field. This survey lays the foundation for developing more capable and reliable AI systems that effectively leverage multimodal dynamic external knowledge bases. Resources are available at https://github.com/llm-lab-org/Multimodal-RAG-Survey.
Ask in Any Modality:
A Comprehensive Survey on Multimodal Retrieval-Augmented Generation
[…]
**footnotetext: These authors contributed equally.
1 Introduction & Background
In recent years, significant breakthroughs have been achieved in language models, driven primarily by the advent of transformers Vaswani et al. (2017), enhanced computational capabilities, and the availability of large-scale training data Naveed et al. (2024). The emergence of foundational Large Language Models (LLMs) Ouyang et al. (2022); Grattafiori et al. (2024); Touvron et al. (2023); Qwen et al. (2025); Anil et al. (2023) has revolutionized natural language processing (NLP), demonstrating unprecedented capabilities in a wide range of tasks including instruction following Qin et al. (2024), sophisticated reasoning Wei et al. (2024), In-context Learning Brown et al. (2020), and multilingual machine translation Zhu et al. (2024a). These advancements have elevated the performance of various NLP tasks, opening new avenues for research and application. Despite their remarkable achievements, LLMs face significant challenges, including hallucination, outdated internal knowledge, and a lack of verifiable reasoning Huang et al. (2024a); Xu et al. (2024b). Their reliance on parametric memory restricts their ability to access up-to-date knowledge, making them less effective for knowledge-intensive tasks compared to task-specific architectures. Moreover, providing provenance for their decisions and updating their world knowledge remain critical open problems Lewis et al. (2020).
Retrieval-Augmented Generation (RAG)
RAG Lewis et al. (2020) has emerged as a promising solution to these limitations by enabling LLMs to retrieve and incorporate external knowledge, improving factual accuracy and reducing hallucinations Shuster et al. (2021); Ding et al. (2024a). By dynamically accessing vast external knowledge repositories, RAG systems enhance knowledge-intensive tasks while ensuring responses remain grounded in verifiable sources Gao et al. (2023). In practice, RAG systems operate through a retriever-generator pipeline. The retriever leverages embedding models Chen et al. (2024b); Rau et al. (2024) to identify relevant passages from external knowledge bases and optionally applies re-ranking techniques to improve retrieval precision Dong et al. (2024a). These retrieved passages are then passed to the generator, incorporating this external context to produce informed responses. Recent advancements in RAG frameworks Asai et al. (2023); An et al. (2024); Lee et al. (2024); Liu et al. (2024c) have introduced iterative inference processes, enabling more accurate and reliable answers by refining both retrieval and generation stages. However, traditional RAG architectures are primarily designed for textual information, limiting their ability to address multimodal challenges that require integrating diverse data formats.
Multimodal Learning
Parallel to these developments, significant advances in multimodal learning have reshaped artificial intelligence by enabling systems to integrate and analyze heterogeneous data sources for a holistic representation of information. The introduction of CLIP (Contrastive Language-Image Pre-training) Radford et al. (2021) marked a pivotal moment in connecting visual and textual information through contrastive learning, inspiring numerous subsequent models and applications Alayrac et al. (2024); Wang et al. (2023); Pramanick et al. (2023). These breakthroughs have driven progress in various domains, including sentiment analysis Das and Singh (2023) and cutting-edge biomedical research Hemker et al. (2024), demonstrating the value of multimodal approaches. By enabling systems to process and understand diverse data types such as text, images, audio, and video, multimodal learning has become essential to advancing artificial general intelligence (AGI) Song et al. (2025).
Multimodal RAG
The extension of LLMs to multimodal LLMs (MLLMs) has further expanded their capabilities, allowing them to process, reason, and generate outputs across diverse modalities Liu et al. (2023a); Team et al. (2024); Li et al. (2023b). For instance, GPT-4 OpenAI et al. (2024) demonstrates human-level performance in various benchmarks by accepting inputs in both text and images, marking a significant milestone in multimodal perception and interaction. Building on this foundation, multimodal RAG systems extend traditional RAG frameworks by incorporating multimodal knowledge sources, such as images and audio, to provide enriched context for generation Hu et al. (2023); Chen et al. (2022a). This integration enhances the precision of generated outputs while leveraging multimodal cues to improve the reasoning capabilities of MLLMs. The general flow of the Multimodal RAG pipeline is illustrated in Figure 1. However, these multimodal systems also present unique challenges, including determining which modalities to retrieve, effectively fusing diverse data types, and addressing the complexities of cross-modal relevance Zhao et al. (2023).
Task Formulation
We present a mathematical formulation of the general task for multimodal RAG systems. These systems generate a multimodal response, denoted as $r$, in response to a query $q$, which is typically in textual format.
Let $D=\{d_{1},d_{2},...,d_{n}\}$ be a corpus consisting of $n$ multimodal documents. Each document $d_{i}$ in $D$ is associated with a modality $M_{d_{i}}$ and is processed by a modality-specific encoder $Enc_{M_{d_{i}}}$:
$z_{i}=Enc_{M_{d_{i}}}(d_{i})$ (1)
The set of all encoded representations is denoted by:
$Z=\{z_{1},z_{2},...,z_{n}\}$ (2)
Modality-specific encoders map different modalities into a shared semantic space for cross-modal alignment. A retrieval model $R$ assesses the relevance of each encoded document representation $z$ with respect to the query $q$, represented as $R(q,z)$. To construct the retrieval-augmented multimodal context, the retrieval model selects the most relevant documents based on a modality-specific threshold:
$X=\{d_{i}\mid s(e_{q},z_{i})\geq\tau_{M_{d_{i}}}\}$ (3)
where $\tau_{M_{d_{i}}}$ is a relevancy threshold for the modality of $M_{d_{i}}$, $e_{q}$ is the encoded representation of $q$ in the shared semantic space, and $s$ is a scoring function that measures the relevance between the encoded query and document representations. The generative model $G$ produces the final multimodal response by conditioning on both the user query $q$ and the retrieved documents $X$ as context:
$r=G(q,X)$ (4)
Related Works
As the field of multimodal RAGs is newly introduced and evolving rapidly, especially in recent years, there is a pressing need for a comprehensive survey that explores the current innovations and frontiers of these systems. While more than ten surveys have been published on RAG-related topics such as Agentic RAG Singh et al. (2025), none provide a detailed and comprehensive overview of advancements in multimodal RAGs. The only related survey to date Zhao et al. (2023) categorizes multimodal RAGs by grouping relevant papers based on their applications and modalities. However, our survey provides a more detailed and different innovation-driven perspective, offering a detailed taxonomy and exploring emerging trends and challenges in depth. Furthermore, significant advancements have been made in the field since its publication, and interest in this topic has grown substantially within the research community. In this survey, we review over 100 papers on multimodal RAGs published in recent years, primarily from the ACL Anthology and other repositories such as the ACM Digital Library.
Contributions
In this work, (i) we provide a comprehensive review of the multimodal RAG field, covering task formulation, datasets, benchmarks, tasks and domain-specific applications, evaluation, and key innovations across retrieval, fusion, augmentation, generation, training strategies, and loss functions. (ii) We introduce a precise structured taxonomy (Figure 2) that categorizes state-of-the-art models based on their primary contributions, highlighting methodological advancements and emerging frontiers. (iii) To support further research, we make resources, including datasets, benchmarks, and key innovations, publicly available. (iv) We identify current research trends and knowledge gaps, providing insights and recommendations to guide future advancements in this evolving field.
Figure 1: Overview of the multimodal retrieval-augmented generation (RAG) pipeline, highlighting the advancements and techniques employed at each stage. The flow begins with query preprocessing, where user queries are refined and then encoded into a shared embedding space alongside a multimodal database. Retrieval strategies, such as modality-centric retrieval, similarity search, and re-ranking, enhance document selection, while fusion mechanisms align and integrate data from multiple modalities using score fusion or attention-based methods. Augmentation techniques such as iterative retrieval with feedback mechanisms, further refine the retrieved documents for the multimodal LLM. The generation stage incorporates innovations like Chain-of-Thought reasoning and source attribution for better outputs, with loss functions combining alignment loss and generation loss to optimize both retrieval and generation components. Noise management techniques are also applied to improve training stability and robustness.
2 Datasets and Benchmarks
Multimodal RAG research employs diverse datasets and benchmarks to evaluate retrieval, integration, and generation across heterogeneous sources. Image–text tasks, including captioning and retrieval, commonly use MS-COCO Lin et al. (2014), Flickr30K Young et al. (2014), and LAION-400M Schuhmann et al. (2021), while visual question answering with external knowledge is supported by OK-VQA Marino et al. (2019) and WebQA Chang et al. (2022). For complex multimodal reasoning, MultimodalQA Talmor et al. (2021) integrates text, images, and tables, whereas video-text tasks leverage ActivityNet Caba Heilbron et al. (2015) and YouCook2 Zhou et al. (2018). In the medical domain, MIMIC-CXR Johnson et al. (2019) and CheXpert Irvin et al. (2019) facilitate tasks such as medical report generation. It is noteworthy that a number of these datasets are unimodal (e.g., solely text-based or image-based). Unimodal datasets are frequently employed to represent a specific modality and are subsequently integrated with complementary datasets from other modalities. This modular approach allows each dataset to contribute its domain-specific strengths, thereby enhancing the overall performance of the multimodal retrieval and generation processes.
Benchmarks assess multimodal RAG systems on visual reasoning, external knowledge integration, and dynamic retrieval. The $M^{2}RAG$ Ma et al. (2024c) benchmark provides a unified evaluation framework that combines fine-grained text-modal and multimodal metrics to jointly assess both the quality of generated language and the effective integration of visual elements Vision-focused evaluations, including MRAG-Bench Hu et al. (2024c), VQAv2 Goyal et al. (2017a) and VisDoMBench Suri et al. (2024), test models on complex visual tasks. Dyn-VQA Li et al. (2024b), MMBench Liu et al. (2025), and ScienceQA Saikh et al. (2022) evaluate dynamic retrieval and multi-hop reasoning across textual, visual, and diagrammatic inputs. knowledge-intensive benchmarks, such as TriviaQA Joshi et al. (2017) and Natural Questions Kwiatkowski et al. (2019), together with document-oriented evaluations such as OmniDocBench Ouyang et al. (2024), measure integration of unstructured and structured data. Advanced retrieval benchmarks such as RAG-Check Mortaheb et al. (2025a) evaluate retrieval relevance and system reliability, while specialized assessments like Counterfactual VQA Niu et al. (2021) test robustness against adversarial inputs. Additionally, OCR impact studies such as OHRBench Zhang et al. (2024d) examine the cascading effects of errors on RAG systems. Additional details about datasets, benchmarks, and their categorization are presented in Table 1 and Table 2 in Appendix B.
3 Evaluation
Assessing multimodal RAG models presents significant challenges due to their diverse modalities, intricate architectures, and the diverse evaluation dimensions involved. Given the broad spectrum of modalities and system components, evaluation requires a combination of metrics drawn from vision-language models (VLMs), generative AI, and retrieval systems. These metrics enable the assessment of various aspects of multimodal RAG, such as text and image generation or information retrieval, either independently or in combination. By analyzing prior works, we identified approximately 60 distinct evaluation metrics used in multimodal RAG studies. More details, including the formulas for the RAG evaluation metrics, can be found in Appendix C. In the following paragraphs, we will examine the most important and widely used metrics for evaluating multimodal RAG.
Retrieval Evaluation
Evaluates how well a system retrieves relevant information from a database or knowledge source. Accuracy, recall, and precision are used to evaluate retrieval and downstream task performance, the F1 score also being used to assess both Recall and Precision simultaneously. A more commonly used metric than recall is Recall@K, which measures the proportion of relevant items appearing within the top K retrieved results. Another important metric Mean Reciprocal Rank (MRR), which is utilized by Adjali et al. (2024); Nguyen et al. (2024).
Modality Evaluation
These evaluations focus on criteria that rely on modalities. Text and image are the most commonly used modalities in multimodal RAG. This category of evaluations considers various aspects of these modalities, including the alignment between generated text and images, the fluency of the text, the quality of image captions, and other related factors. In text similarity and overlap metrics, Exact Match (EM), BLEU Papineni et al. (2002), ROUGE Lin (2004) and METEOR Banerjee and Lavie (2005) are used. MultiRAGen Shohan et al. (2024) uses Multilingual ROUGE for multilingual settings. For image captioning evaluation, CIDEr (Consensus-Based Image Description Evaluation) Vedantam et al. (2015) is a widely used metric for evaluating image captioning quality. It quantifies the alignment between a generated caption and a set of human-written reference captions by leveraging the term frequency-inverse document frequency (TF-IDF) and cosine similarity. Yasunaga et al. (2023); Zhao et al. (2024); Luo et al. (2024a); Yuan et al. (2024); Sharifymoghaddam et al. (2024); Hu et al. (2023); Rao et al. (2024); Xu et al. (2024a); Kim et al. (2024); Zhang et al. (2024c) use CIDEr for evaluation. SPICE (Semantic Propositional Image Caption Evaluation) Anderson et al. (2016) is a metric that emphasizes semantic aspects, in contrast to CIDEr. SPIDEr Liu et al. (2017), utilized in Zhang et al. (2024c), is a linear combination of both SPICE and CIDEr. In semantic alignment evaluations, BERTScore Zhang et al. (2020), employed in Sun et al. (2024b); Shohan et al. (2024), measures semantic similarity by comparing BERT-generated embeddings. Chen et al. (2022a); Zhi Lim et al. (2024); Ma et al. (2024c) evaluate Fluency by employing BERTScore. CLIP Score Hessel et al. (2021), used in Sharifymoghaddam et al. (2024); Zhang et al. (2024c), measures image-text similarity using the CLIP model Radford et al. (2021). FID (Fréchet Inception Distance) Heusel et al. (2017) quantifies image quality by computing the distributional difference between the generated and original image features. Yasunaga et al. (2023); Zhao et al. (2024); Sharifymoghaddam et al. (2024); Zhang et al. (2024c) use FID in their works. Zhao et al. (2024) use Kernel Inception Distance (KID) Bińkowski et al. (2018) for evaluation. KID is another metric used to evaluate the quality of generated images. It is similar to FID but offers an unbiased estimator. Zhi Lim et al. (2024) employs the Inception Score (IS) to assess the diversity and quality of generated images based on classification probabilities obtained from the Inception network. For the evaluation of audio synthesis, Zhang et al. (2024c) assesses overall sound quality (OVL) and relevance to the text (REL) with the assistance of human annotators. Additionally, this study employs the Fréchet Audio Distance (FAD) Kilgour et al. (2019) as an evaluation metric. FAD serves a similar purpose to the FID but is specifically designed for the audio domain.
FLOPs, execution time, response time, and average retrieval time per query are utilized to assess efficiency and computational performance Nguyen et al. (2024); Strand et al. (2024); Dang (2024); Zhou (2024). Zhong et al. (2024) employs Spearman’s Rank Correlation (SRC), which calculates the Pearson correlation between the rank values of two variables. Several evaluation metrics are domain- and problem-specific. Geodesic distance measurement quantifies the model’s accuracy by calculating the geodesic distance between predicted and actual geographical coordinates Zhou et al. (2024e). Lahiri and Hu (2024) utilizes Clinical Relevance (CR) for evaluation.
4 Key Innovations and Methodologies
4.1 Retrieval Strategy
Efficient Search and Similarity Retrieval
Modern multimodal RAG systems first encode diverse modalities (text, images, tables) into a unified embedding space to enable direct cross-modal retrieval. The foundational CLIP-Radford et al. (2021) or BLIP-based Li et al. (2022a) approaches leverage contrastive learning, which aligns text and images into a shared latent space via dual-tower encoders. Recent advancements extend this paradigm with novel multimodal retrieval architectures and training strategies Zhou et al. (2024c); Wei et al. (2025); Zhang et al. (2024j). Multimodal RAG systems rely on efficient search strategies to retrieve relevant external knowledge from large-scale multimodal memory. Since dual-encoder models map multimodal inputs into a shared high-dimensional embedding space, maximum inner product search (MIPS) is commonly used for fast and direct similarity comparisons. Systems such as MuRAG Chen et al. (2022a) and RA-CM3 Yasunaga et al. (2023) employ approximate MIPS to efficiently retrieve the top-k candidates by maximizing the inner product between the query vector and a large collection of image–text embeddings. Large-scale implementations leverage distributed MIPS techniques, such as TPU-KNN Chern et al. (2022), to enable high-speed retrieval across extensive datasets. Additionally, ScaNN (Scalable Nearest Neighbors) Guo et al. (2020), MAXSIM score Chan and Ng (2008); Cho et al. (2024), and approximate KNN methods Caffagni et al. (2024) have been adopted for efficient image-text similarity computation.
Recent advancements in MIPS optimization focus on reducing retrieval latency and improving accuracy. These include adaptive quantization Li et al. (2024a); Huang et al. (2024b), hybrid sparse-dense representations Nguyen et al. (2024); Ma et al. (2024a); Zhang et al. (2024a), and learned index structures Zhai et al. (2023); Basnet et al. (2024). These techniques help optimize retrieval performance by balancing efficiency and precision in multimodal RAG systems.
Modality-Based Retrieval
Modality-aware retrieval techniques optimize retrieval efficiency by leveraging the unique characteristics of each modality. Text-centric retrieval remains foundational in multimodal RAG systems. Traditional methods like BM25 Robertson and Zaragoza (2009), along with dense retrievers such as BGE-1.5 Xiao et al. (2023) and BGE-M3 Chen et al. (2024a) dominate text-based evidence retrieval Chen et al. (2022b); Suri et al. (2024); Nan et al. (2024). Novel approaches also address the need for fine-grained semantic matching and domain specificity: for instance, ColBERT Khattab and Zaharia (2020) and PreFLMR Lin et al. (2024b) employ token-level interaction mechanisms that preserve nuanced textual details to improve precision for multimodal queries, while RAFT Zhang et al. (2024i) and CRAG Yan et al. (2024) enhance retrieval performance by ensuring precise citation of text spans, thereby optimizing domain-specific question answering within multimodal RAG pipelines. Vision-centric retrieval focuses on directly leveraging image representations for knowledge extraction Kumar and Marttinen (2024); Yuan et al. (2023). Systems such as EchoSight Yan and Xie (2024) and ImgRet Shohan et al. (2024) retrieve visually similar content by using reference images as queries. Composed Image Retrieval (CMI) models Feng et al. (2023); Zhao et al. (2024); Jang et al. (2024) enhance retrieval by integrating multiple image features into a unified query representation. Similarly, Pic2word Saito et al. (2023) maps visual content to textual descriptions, enabling zero-shot image retrieval. Video-centric retrieval extends vision-based techniques by incorporating temporal dynamics and large video-language models (LVLMs), driven by novel frameworks like iRAG Arefeen et al. (2024), which introduces incremental retrieval for sequential video understanding, and MV-Adapter Jin et al. (2024), optimizing multimodal transfer learning for video-text retrieval. Recent breakthroughs focus on long-context processing: Video-RAG Luo et al. (2024b) leverages visually aligned auxiliary texts (OCR/ASR) to enhance retrieval without proprietary models, while VideoRAG Ren et al. (2025) employs dual-channel architectures with graph-based knowledge grounding for extreme-length videos. For temporal reasoning, CTCH Shen et al. (2024) uses contrastive transformer hashing to model long-term dependencies, and RTime Du et al. (2024) introduces reversed-video hard negatives to rigorously benchmark temporal causality. Meanwhile, for complex video understanding, OmAgent Zhang et al. (2024e) adopts a divide-and-conquer framework, while DRVideo Ma et al. (2024d) addresses long-video understanding by a document-based retrieval approach.
Document Retrieval and Layout Understanding
Recent research has advanced beyond traditional uni-modal retrieval methods toward models that directly process entire document pages. These approaches integrate textual, visual, and structural elements like tables, images, font styles, and page layouts to enhance retrieval performance in complex documents. ColPali Faysse et al. (2024) pioneers end-to-end document image retrieval by embedding page patches with a vision-language backbone, bypassing OCR entirely. Models like ColQwen2 Wang et al. (2024b); Khattab and Zaharia (2020) and M3DocVQA Cho et al. (2024) extend this paradigm with dynamic resolution handling and holistic multi-page reasoning. Newer frameworks refine efficiency and layout understanding: ViTLP Mao et al. (2024) and DocLLM Wang et al. (2024a) pre-train generative models to align spatial layouts with text, while CREAM Zhang et al. (2024b) employs coarse-to-fine retrieval with multimodal efficient tuning to balance accuracy and computation costs. Finally, mPLUG-DocOwl 1.5 Hu et al. (2024a) and 2 Hu et al. (2024b) unify structure learning across formats (e.g., invoices, forms) without OCR dependencies. Together, these approaches highlight a shift toward holistic, layout-sensitive retrieval that leverages visual semantics and generative pre-training.
Re-ranking and Selection Strategies
Effective retrieval in multimodal RAG systems requires not only identifying relevant information but also refining and prioritizing retrieved candidates. Re-ranking and selection strategies enhance retrieval quality by optimizing example selection, refining relevance scoring, and applying filtering mechanisms. Optimized example selection techniques often employ multi-step retrieval, integrating both supervised and unsupervised selection strategies Luo et al. (2024a). Su et al. (2024a) refine multimodal inputs using probabilistic control keywords to improve credibility, while RULE Xia et al. (2024b) calibrates retrieved context selection through statistical methods like the Bonferroni correction to mitigate factuality risks. Furthermore, selective fusion mechanisms dynamically refine the retrieval outcomes Yuan et al. (2023), and key-frame extraction techniques ensure diversity in video-based retrieval through clustering-based frame selection Dong et al. (2024b). Several methods employ advanced scoring mechanisms to improve retrieval relevance Mortaheb et al. (2025b, a). Multimodal similarity measures, including negative log-likelihood (NLL), structural similarity index measure (SSIM) Wang et al. (2020), normalized cross-correlation (NCC), and BERTScore Zhang et al. (2020), aid in re-ranking retrieved documents based on textual and visual similarity. Cross-encoders trained on sequence classification tasks refine relevance scoring within retrieval pipelines Zhi Lim et al. (2024), while hierarchical post-processing techniques integrate retrieval, passage-level, and answer confidence scores for improved ranking Zhang et al. (2024h). LDRE Yang et al. (2024) employs semantic ensemble methods to adaptively weigh multiple caption features, whereas RAGTrans Cheng et al. (2024) and OMG-QA Nan et al. (2024) integrate traditional ranking functions like BM25 Robertson and Zaragoza (2009). Advanced re-ranking techniques also incorporate Q-Former-based image-to-text correspondence for fine-grained retrieval Yan and Xie (2024); Xu et al. (2024a). Advanced filtering methods are crucial for enhancing multi-modal RAG performance by ensuring that only high-quality, relevant data is processed. Hard negative mining techniques, as seen in GME Zhang et al. (2024j) and MM-Embed Lin et al. (2024a), use modality-aware sampling and synthesized negatives from imbalanced data to mitigate modality bias and refine contrastive learning. Similarly, consensus-based filtering approaches, exemplified by MuRAR Zhu et al. (2024d) and ColPali Faysse et al. (2024), employ source attribution and multi-vector mapping to filter out low-similarity candidates and reduce hallucinations in enterprise QA. Moreover, dynamic modality filtering methods—including RAFT Zhang et al. (2024i), Img2Loc Zhou et al. (2024e), and MAIN-RAG Chang et al. (2024)—train retrievers to disregard confusing data, thereby enhancing the discriminative capacity and overall robustness of multi-modal retrieval systems.
4.2 Fusion Mechanisms
Score Fusion and Alignment
Recent multimodal RAG models emphasize aligning embeddings or fusing scores across different modalities to enable adaptive retrieval. Zhi Lim et al. (2024) transform text, tables, and images into a unified textual representation using a cross-encoder trained for relevance scoring. To address multimodal LLM input constraints, Sharifymoghaddam et al. (2024) introduce interleaved image-text pairs, merging multiple few-shot images vertically for compatibility with models like LLaVA Liu et al. (2023a) while aligning modalities via CLIP score fusion Hessel et al. (2021) and BLIP feature fusionLi et al. (2022a). For fine-grained alignment, REVEAL Hu et al. (2023) injects retrieval scores into attention layers and minimizes L2-Norm distances between query and knowledge embeddings, and MA-LMM He et al. (2024) aligns video-text embeddings via a BLIP-inspired Query Transformer (Q-Former)Li et al. (2022a). In contrast, LLM-RA Jian et al. (2024) concatenates text and visual embeddings into joint queries to reduce retrieval noise. RA-BLIP Ding et al. (2024b) employs a 3-layer BERT-based adaptive fusion module to unify visual-textual semantics. CLIP-based alignment strategies are widely adopted; Riedler and Langer (2024) embed images and questions into a shared CLIP space, while VISA Ma et al. (2024b) uses the Document Screenshot Embedding (DSE) model as the retrieval component of the RAG system to align and fuse textual queries with visual document representations. DSE directly encodes both textual queries and document screenshots into a shared embedding space, allowing efficient retrieval based on cosine similarity. Xue et al. (2024) map object-predicate pairs into a shared semantic space using a prototype-based embedding network (PENET) Zheng et al. (2023), aligning visual features with textual prototypes. Re-IMAGEN Chen et al. (2022b) refines text-to-image synthesis with interleaved classifier-free guidance during diffusion sampling, balancing creativity and entity fidelity. To improve multimodal alignment, MegaPairs Zhou et al. (2024a) integrates CLIP-based and MLLM-based retrieval through score fusion. Meanwhile, Wiki-LLaVA Caffagni et al. (2024) employs CLIP encoders for hierarchical document retrieval. VISRAG Yu et al. (2024) refines alignment via position-weighted mean pooling on VLM hidden states, prioritizing later tokens for relevance. RAG-Driver Yuan et al. (2024) aligns visual-language embeddings via visual instruction tuning and an MLP projector.
Attention-Based Mechanisms
Attention-based fusion dynamically weights cross-modal interactions for task-specific reasoning. RAMM Yuan et al. (2023) employs a dual-stream co-attention transformer, combining self-attention (intra-modal) and cross-attention (inter-modal) to fuse retrieved biomedical images/texts with input data. Similarly, EMERGE Zhu et al. (2024b) integrates Electronic Health Records (EHR) data with a cross-attention adaptive network, while MORE Cui et al. (2024) uses cross-attention to weigh multimodal results for commonsense reasoning. For user-centric tasks, RAGTrans Cheng et al. (2024) integrates user-aware attention to fuse social media features, and AlzheimerRAG Lahiri and Hu (2024) applies cross-modal attention to synthesize text, images, and tables. Video-text alignment is advanced by MV-Adapter Jin et al. (2024), which leverages a Cross Modality Tying (CMT) to align embeddings, and M2-RAAP Dong et al. (2024b), enhances video-text fusion using an Auxiliary-Caption-Guided (ACG) strategy. It employs Frame-Caption Re-weighting (FCR) to boost frames with high inter-modal consistency and Text-Caption Re-weighting (TCR) to adjust weights based on intra-modal similarity. A Mutual-guided (Mug) Alignment Head then filters misaligned features via a dot-product similarity and frame-to-token attention, producing refined frame-specific text representations. Generative refinement is demonstrated in Xu et al. (2024a), where gated cross-attention conditions text generation on visual features. Similarly, C3Net Zhang et al. (2024c) enhances alignment between queries and candidates within CLIP’s contrastive space. Mu-RAG Chen et al. (2022a) uses intermediate fusion with cross-attention for open QA, balancing modality-specific and joint representations. Kim et al. (2024) leverage cross-modal memory retrieval with pre-trained CLIP ViT-L/14 to map video-text pairs into a shared space, enabling dense captioning through the attention-based fusion of retrieved memories.
Unified Frameworks and Projections
Architectures unifying multimodal inputs through frameworks or projections address scalability. Su et al. (2024a) employ hierarchical cross-chains and late fusion for healthcare data, while IRAMIG Liu et al. (2024b) iteratively consolidates multimodal results into knowledge representations. For document understanding, M3DocRAG Cho et al. (2024) flattens multi-page documents into a unified embedding tensor, and PDF-MVQA Ding et al. (2024c) combines Region-of-Interest (RoI)-based (VisualBERT) and patch-based (CLIP) VLPMs Long et al. (2022). DQU-CIR Wen et al. (2024) exemplifies raw-data unification by converting images into text captions for complex queries and overlaying text onto images for simple queries. The model then fuses embeddings using MLP-learned weights. SAM-RAG Zhai (2024) aligns image-text modalities by generating captions for images, converting the multimodal input into unimodal text representation for subsequent processing. UFineBench Zuo et al. (2024) uses a shared granularity decoder for ultra-fine text-person retrieval. Nguyen et al. (2024) introduce Dense2Sparse (D2S) projection, converting dense embeddings from models like BLIP/ALBEF Li et al. (2022a) into sparse lexical vectors via layer normalization and probabilistic expansion control, optimizing storage and interpretability.
4.3 Augmentation Techniques
Basic RAG systems typically follow a single retrieval step and pass retrieved content directly to the generation phase without further refinement, which can lead to inefficiencies and suboptimal outputs. To address this, augmentation techniques have been introduced to enhance the utility of retrieved data before generation, ensuring that the model can better interpret, structure, and integrate multimodal information Gao et al. (2023).
Context Enrichment
This focuses on enhancing the relevance of retrieved knowledge for the generation phase by refining external knowledge or retrieved data. General approaches such as EMERGE Zhu et al. (2024b) integrate entity relationships and semantic descriptions to provide richer context. MiRAG Adjali et al. (2024) improves retrieval effectiveness by expanding initial queries through entity retrieval and query reformulation, enhancing subsequent stages for the visual question-answering task. Wiki-LLaVA Caffagni et al. (2024) incorporates text chunks and image tokens into multimodal inputs to enrich prompts for generation. Video-RAG Luo et al. (2024b) enhances long-video understanding through Query Decoupling, which reformulates user queries into structured retrieval requests to extract additional auxiliary multimodal context. Img2Loc Zhou et al. (2024e) enhances accuracy by including both the most similar and the most dissimilar points from the database in the prompt, allowing the model to rule out implausible locations for its predictions. Finally, Xue et al. (2024) introduce semantic-enhanced prompts, aligning structured image data with queries for contextual accuracy.
Adaptive and Iterative Retrieval
For more complex queries, dynamic retrieval mechanisms have proven effective. SKURG Yang et al. (2023) employs adaptive retrieval steps to dynamically adjust the number of hops based on the query complexity. IRAMIG Liu et al. (2024b) iteratively refines multimodal queries by incorporating feedback from previous iterations and context created by the fusion of retrieved items from text, images, and audio modalities. Similarly, OMG-QA Nan et al. (2024) introduces a multi-round retrieval strategy, where each retrieval step incorporates episodic memory to refine subsequent queries. An evaluation module assesses retrieval effectiveness at each step, guiding the refinement of subsequent retrieval efforts through feedback. Taking a different approach, SAM-RAG Zhai (2024) further improves relevance by dynamically selecting the optimal number of retrieved contexts for each query. MMed-RAG Xia et al. (2024a) uses context filtering to discard low-relevance results. OmniSearch Li et al. (2024b) introduces a self-adaptive retrieval agent that dynamically decomposes complex multimodal questions into sub-question chains and plans retrieval actions in real-time based on retrieved content. mR2AG Zhang et al. (2024g) introduces reflection mechanisms to enhance retrieval effectiveness. Retrieval-Reflection dynamically determines when external knowledge is necessary, while Relevance-Reflection filters retrieved content, retaining only task-critical evidence. Similarly, RAGAR Khaliq et al. (2024) iteratively refines retrieved data by leveraging prior responses and multimodal analysis, ensuring contextual consistency.
4.4 Generation Techniques
Recent advancements in multimodal RAG generation focus on robustness, cross-modal coherence, and task-specific adaptability. These innovations can be broadly categorized into these sections:
In-Context Learning
In-context learning (ICL) with retrieval augmentation enhances reasoning in multimodal RAGs by leveraging retrieved content as few-shot examples without requiring retraining. Models such as RMR Tan et al. (2024), Sharifymoghaddam et al. (2024), and RA-CM3 Yasunaga et al. (2023), extend this paradigm to multimodal RAG settings. RAG-Driver Yuan et al. (2024) refines ICL by retrieving relevant driving experiences from a memory database, ensuring scenario-specific contextual alignment. MSIER Luo et al. (2024a) further enhances example selection through a Multimodal Supervised In-Context Examples Retrieval framework, leveraging a foundation MLLM scorer to evaluate both textual and visual relevance. Meanwhile, Raven Rao et al. (2024) introduces Fusion-in-Context Learning, a novel approach that enriches ICL by incorporating a more diverse set of in-context examples, leading to better performance than standard ICL.
Reasoning
Structured reasoning techniques, such as chain-of-thought (CoT), decompose complex reasoning into smaller sequential steps, enhancing coherence and robustness in multimodal RAG systems. RAGAR Khaliq et al. (2024) introduces Chain of RAG and Tree of RAG to iteratively refine fact-checking queries and explore branching reasoning paths for more robust evidence generation. VisDoM Suri et al. (2024) integrates CoT with evidence curation to ensure logical and contextual accuracy, while SAM-RAG Zhai (2024) employs reasoning chains alongside multi-stage answer verification to enhance the relevance, utility, and support of generated responses. Meanwhile, LDRE Yang et al. (2024) leverages LLMs for divergent compositional reasoning, generating refined captions by incorporating dense captions and modification text.
Instruction Tuning
Several works have fine-tuned generation components for specific applications. RA-BLIP Ding et al. (2024b), uses Q-Former architecture from InstructBLIP Dai et al. (2023) to extract visual features based on question instructions. RAGPT Lang et al. (2025) uses a context-aware prompter, which captures contextual knowledge from relevant instances and generates dynamic prompts. mR2AG Zhang et al. (2024g) uses instruction tuning with the mR2AG-IT dataset to train MLLMs to adaptively invoke retrieval, identify relevant evidence, and generate accurate answers for knowledge-based VQA tasks. RagVL Chen et al. (2024d) employs instruction tuning to enhance the ranking capability of MLLMs, serving them as a re-ranker for filtering the top-k retrieved images. Jang et al. (2024) propose a Visual Delta Generator to discriminate differences between image pairs and produce a textual response. MMed-RAG Xia et al. (2024a) uses preference fine-tuning to help the model learn when to rely on retrieved information and when to trust its internal knowledge. Additionally, several methods have been proposed to improve generation quality by constructing datasets from the previous mistakes of LLMs and fine-tuning them accordingly. MegaPairs Zhou et al. (2024a) synthesizes large-scale multimodal instruction-tuning datasets for this purpose. Similarly, Surf Sun et al. (2024a) creates a dataset of positive and negative pairs derived from the prior errors of LVLMs and subsequently fine-tunes the LVLM using this dataset. Additionally, Rule Xia et al. (2024b) refines Med-LVLM by applying direct preference optimization on curated preference data, specifically focusing on instances where overreliance on retrieved contexts resulted in errors.
Source Attribution and Evidence Transparency
Ensuring source attribution in multimodal RAG systems is a key focus of recent research. MuRAR Zhu et al. (2024d) integrates multimodal data, fetched by a source-based retriever, to refine LLM’s initial response, ensuring coherence and informativeness. VISA Ma et al. (2024b) uses large vision-language models (VLMs) to generate answers with visual source attribution by identifying and highlighting supporting evidence in retrieved document screenshots. Similarly, OMG-QA Nan et al. (2024) enhances transparency by prompting the LLM to explicitly cite evidence in generated responses.
4.5 Training Strategies
Training multimodal RAG models involves a multi-stage process to effectively handle cross-modal interactions Chen et al. (2022a). Pretraining establishes the foundation using large paired datasets to learn cross-modal relationships while fine-tuning adapts models to specific tasks by leveraging cross-modal attention Ye et al. (2019). For instance, REVEAL Hu et al. (2023) integrates multiple training objectives. Its pretraining phase optimizes Prefix Language Modeling Loss ($L_{\text{PrefixLM}}$), where the model predicts text continuations from a prefix and an image. Supporting losses include Contrastive Loss ($L_{\text{contra}}$) for aligning queries with pseudo-ground-truth knowledge, Disentangled Regularization Loss ($L_{\text{decor}}$) to improve embedding expressiveness, and Alignment Regularization Loss ($L_{\text{align}}$) to align query and knowledge embeddings. During fine-tuning, a cross-entropy objective trains the model for tasks like VQA and image captioning.
Alignment
Contrastive learning improves representation quality by pulling positive pairs closer and pushing negative pairs apart in the embedding space. A common objective is the InfoNCE loss van den Oord et al. (2019), which maximizes the mutual information between positive pairs while minimizing similarity to negatives. Several multimodal RAG models, such as VISRAG Yu et al. (2024), MegaPairs Zhou et al. (2024a) and SAM-RAG Zhai (2024) utilize InfoNCE loss to improve retrieval-augmented generation. Furthermore, EchoSight Yan and Xie (2024) enhances retrieval accuracy by selecting visually similar yet contextually distinct negatives, while HACL Jiang et al. (2024) improves generation by introducing hallucinative captions as distractors. Similarly, UniRaG Zhi Lim et al. (2024) strengthens retrieval by incorporating hard negative documents, helping the model distinguish relevant contexts from noise.
The eCLIP loss Kumar and Marttinen (2024) extends contrastive training by integrating expert-annotated data and an auxiliary Mean Squared Error (MSE) loss to refine embedding quality. Mixup strategies further augment training by generating synthetic positive pairs, improving generalization in contrastive learning Kumar and Marttinen (2024). Dense2Sparse Nguyen et al. (2024) incorporates two unidirectional losses: an image-to-caption loss $\ell(I\to C)$ and a caption-to-image loss $\ell(C\to I)$. It enforces sparsity through L1 regularization, optimizing retrieval precision by balancing dense and sparse representations. Details of formulas for widely used RAG loss functions can be found in Appendix D.
Robustness
Multimodal training faces challenges such as noise and modality-specific biases Buettner and Kovashka (2024). Managing noisy retrieval inputs is critical for maintaining model performance. MORE Cui et al. (2024) injects irrelevant results during training to enhance focus on relevant inputs. AlzheimerRAG Lahiri and Hu (2024) uses progressive knowledge distillation to reduce noise while maintaining multimodal alignment. RAGTrans Cheng et al. (2024) leverages hypergraph-based knowledge aggregation to refine multimodal representations, ensuring more effective propagation of relevant information. RA-BLIP Ding et al. (2024b) introduces the Adaptive Selection Knowledge Generation (ASKG) strategy, which leverages the implicit capabilities of LLMs to filter relevant knowledge for generation through a denoising-enhanced loss term, eliminating the need for fine-tuning. This approach achieves strong performance compared to baselines while significantly reducing computational overhead by minimizing trainable parameters. RagVL Chen et al. (2024d) improves robustness through noise-injected training by adding hard negative samples at the data level and applying Gaussian noise with loss reweighting at the token level, enhancing the model’s resilience to multimodal noise. Finally, RA-CM3 Yasunaga et al. (2023) enhances generalization using Query Dropout, which randomly removes query tokens during retrieval, serving as a regularization method that improves generator performance.
{forest}
forked edges, for tree= child anchor=west, parent anchor=east, grow’=east, anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, minimum height=2em, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0.4em, inner ysep=0.6em, line width=0.8pt, text width=8.5em, ver/.style= fill=gray!10, rotate=90, child anchor=north, parent anchor=south, anchor=center, text width=8em , leaf/.style= text opacity=1, inner sep=2pt, fill opacity=.5, fill=hidden-blue!90, text=black, text width=44.5em font=, inner xsep=0.4em, inner ysep=0.6em, draw, , , [Multimodal RAG, ver [ Retrieval
Strategies (§4.1) [ Efficient-Search and
Similarity Retrieval
 (§4.1) [ Maximum Inner Product Search (MIPS) [ TPU-KNN Chern et al. (2022), ScaNN Guo et al. (2020), MAXSIM score Chan and Ng (2008), OmniSearch Li et al. (2024b), MuRAG Chen et al. (2022a), RetrievalAttention Liu et al. (2024a), RA-CM3 Yasunaga et al. (2023), RAVEN Rao et al. (2024) FactMM-RAG Sun et al. (2024b) , leaf, text width=44.5em ] ] [ Multimodal
Encoders [ CLIP Radford et al. (2021), BLIP Li et al. (2022a), MARVEL Zhou et al. (2024c), ALIGN Jia et al. (2021), FLAVA Singh et al. (2022), UniVL-DR Liu et al. (2023b), UniIR Wei et al. (2025), GME Zhang et al. (2024j), VISTA Zhou et al. (2024b), ColPali Faysse et al. (2024), InternVideo Wang et al. (2022), Ovis Lu et al. (2024) Mi-RAG Adjali et al. (2024) , leaf, text width=44.5em ] ] ] [ Modality-Centric
Retrieval (§4.1) [ Text-Centric [ Contriever Izacard et al. (2022), GTE Li et al. (2023c), Re-Imagen Chen et al. (2022b), BM25 Robertson and Zaragoza (2009), BGE-1.5 Xiao et al. (2023), BGE-M3 Chen et al. (2024a), CapRet Shohan et al. (2024), OMG-QA Nan et al. (2024), ColBERT Khattab and Zaharia (2020), PreFLMR Lin et al. (2024b), RAFT Zhang et al. (2024i), CRAG Yan et al. (2024), M2-RAG Ma et al. (2024c) , leaf, text width=44.5em ] ] [ Vision-Centric [ VQA4CIR Feng et al. (2023), Unifashion Zhao et al. (2024), Jang et al. Jang et al. (2024), Pic2word Saito et al. (2023), eClip Kumar and Marttinen (2024), RAMM Yuan et al. (2023), Joshi et al. Joshi et al. (2024), VISA Ma et al. (2024b), ImgRet Shohan et al. (2024), EchoSight Yan and Xie (2024), Xue et al. Xue et al. (2024) , leaf, text width=44.5em ] ] [ Video-Centric [ iRAG Arefeen et al. (2024), VideoRAG Ren et al. (2025), VideoRAG Jeong et al. (2025), MV-Adapter Jin et al. (2024), OmAgent Zhang et al. (2024f), CM2 Kim et al. (2024), Video-RAG Luo et al. (2024b), CTCH Shen et al. (2024), RTime Du et al. (2024), VideoMAE Tong et al. (2022), DrVideo Ma et al. (2024d) , leaf, text width=44.5em ] ] [ Document Retrieval [ ColPali Faysse et al. (2024), ColQwen2 Wang et al. (2024b), M3DocVQA Cho et al. (2024), ViTLP Mao et al. (2024), DocLLM Wang et al. (2024a), CREAM Zhang et al. (2024b), mPLUG-DocOwl 1.5 Hu et al. (2024a), mPLUG-DocOwl 2 Hu et al. (2024b), VisDom Suri et al. (2024), DSE Ma et al. (2024a) , leaf, text width=44.5em ] ] ] [ Re-ranking
Strategies (§4.1) [ Optimized Example
Selection [ MSIER Luo et al. (2024a), Hybrid RAG Su et al. (2024b), RULE Xia et al. (2024b), RAMM Yuan et al. (2023), M2RAAP Dong et al. (2024b) , leaf, text width=44.5em ] ] [ Relevance Score
Evaluation [ RAG-Check Mortaheb et al. (2025a, b), UniRaG Zhi Lim et al. (2024), MR2AG Zhang et al. (2024h), LDRE Yang et al. (2024), BM25 Robertson and Zaragoza (2009), RAGTrans Cheng et al. (2024), OMG-QA Nan et al. (2024), EchoSight Yan and Xie (2024), EgoInstructor Xu et al. (2024a) , leaf, text width=44.5em ] ] [ Filtering
Mechanisms [ MAIN-RAG Chang et al. (2024), MM-Embed Lin et al. (2024a), GME Zhang et al. (2024j), Img2Loc Zhou (2024), MuRAR Zhu et al. (2024d) RAFT Zhang et al. (2024i) , leaf, text width=44.5em ] ] ] ] [ Fusion
Mechanisms (§4.2) [ Score Fusion
and Alignment
 (§4.2) [ M3 Cai et al. (2025)  Zhi Lim et al. (2024),  Sharifymoghaddam et al. (2024), REVEAL Hu et al. (2023), RAG-Driver Yuan et al. (2024), LLM-RA Jian et al. (2024),  Riedler and Langer (2024), VISA Ma et al. (2024b), MA-LMM He et al. (2024),  Xue et al. (2024), RA-BLIP Ding et al. (2024b), Re-IMAGEN Chen et al. (2022b), MegaPairs Zhou et al. (2024a), Wiki-LLaVA Caffagni et al. (2024), VISRAG Yu et al. (2024) , leaf, text width=55.0em ] ] [ Attention-Based
Mechanisms (§4.2) [ RAMM Yuan et al. (2023), EMERGE Zhu et al. (2024b), MORE Cui et al. (2024), RAGTrans Cheng et al. (2024), AlzheimerRAG Lahiri and Hu (2024), MV-Adapter Jin et al. (2024), C3Net Zhang et al. (2024c),  Xu et al. (2024a),  Kim et al. (2024), M2-RAAP Dong et al. (2024b), Mu-RAG Chen et al. (2022a), Ou et al. Ou et al. (2025), CADMR Khalafaoui et al. (2024) , leaf, text width=55.0em ] ] [ Unified Frameworks
and Projections
 (§4.2) [ Hybrid-RAG Su et al. (2024b), Dense2Sparse Nguyen et al. (2024), IRAMIG Liu et al. (2024b), M3DocRAG Cho et al. (2024), DQU-CIR Wen et al. (2024), PDF-MVQA Ding et al. (2024c), SAM-RAG Zhai (2024), UFineBench Zuo et al. (2024),  Li et al. (2022a), , leaf, text width=55.0em ] ] ] [ Augmentation
Techniques (§4.3) [ Context-Enrichment (§4.3) [ EMERGE Zhu et al. (2024b), MiRAG Adjali et al. (2024), Wiki-LLaVA Caffagni et al. (2024), Video-RAG Luo et al. (2024b), Img2Loc Zhou et al. (2024e),  Xue et al. (2024) , leaf, text width=55.0em ] ] [ Adaptive and
Iterative
Retrieval (§4.3) [ SKURG Yang et al. (2023), IRAMIG Liu et al. (2024b), OMG-QA Nan et al. (2024), SAM-RAG Zhai (2024), MMed-RAG Xia et al. (2024a), OmniSearch Li et al. (2024b), mR2AG Zhang et al. (2024g), RAGAR Khaliq et al. (2024) , leaf, text width=55.0em ] ] ] [ Generation
Techniques (§4.4) [ In-Context
Learning (§4.4) [ RMR Tan et al. (2024),  Sharifymoghaddam et al. (2024), RA-CM3 Yasunaga et al. (2023), RAG-Driver Yuan et al. (2024), MSIER Luo et al. (2024a), Raven Rao et al. (2024) , leaf, text width=55.0em ] ] [ Reasoning (§4.4) [ RAGAR Khaliq et al. (2024), VisDoM Suri et al. (2024), SAM-RAG Zhai (2024), LDRE Yang et al. (2024) , leaf, text width=55.0em ] ] [ Instruction
Tuning (§4.4) [ RA-BLIP Ding et al. (2024b), RAGPT Lang et al. (2025), mR2AG Zhang et al. (2024g), RagVL Chen et al. (2024d), Jang et al. (2024), MMed-RAG Xia et al. (2024a), MegaPairs Zhou et al. (2024a), Surf Sun et al. (2024a), Rule Xia et al. (2024b) , leaf, text width=55.0em ] ] [ Source
Attribution (§4.4) [ MuRAR Zhu et al. (2024d), VISA Ma et al. (2024b), OMG-QA Nan et al. (2024) , leaf, text width=55.0em ] ] ] [ Training Strategies (§4.5) [ Alignment (§4.5) [ VISRAG Yu et al. (2024), MegaPairs Zhou et al. (2024a), SAM-RAG Zhai (2024), EchoSight Yan and Xie (2024), HACL Jiang et al. (2024),  Zhi Lim et al. (2024),  Kumar and Marttinen (2024), Dense2Sparse Nguyen et al. (2024) , leaf, text width=55.0em ] ] [ Robustness (§4.5) [  Buettner and Kovashka (2024), MORE Cui et al. (2024), AlzheimerRAG Lahiri and Hu (2024), RAGTrans Cheng et al. (2024), RA-BLIP Ding et al. (2024b), RagVL Chen et al. (2024d), RA-CM3 Yasunaga et al. (2023) , leaf, text width=55.0em ] ] ] ] ]
Figure 2: Taxonomy of recent advances and enhancements in multimodal retrieval-augmented generation research.
5 Tasks Addressed by Multimodal RAGs
Multimodal RAG systems extend RAG beyond unimodal settings to tasks requiring cross-modal integration, enhancing performance across modalities such as text, images, and audio. In content generation, these models enhance image captioning Zhi Lim et al. (2024); Hu et al. (2023); Rao et al. (2024) and text-to-image synthesis Yasunaga et al. (2023); Chen et al. (2022b) by retrieving relevant contextual information. They also improve coherence in visual storytelling and ensure factual alignment in multimodal summarization Tonmoy et al. (2024). In knowledge-intensive applications, multimodal RAG supports open-domain and knowledge-seeking question answering Chen et al. (2024d); Ding et al. (2024b); Yuan et al. (2023), video-based QA Luo et al. (2024b), and fact verification Khaliq et al. (2024), grounding responses in retrieved knowledge and thereby mitigating hallucinations. Cross-modal retrieval further advances zero-shot image–text retrieval Yang et al. (2024); Dong et al. (2024b). Additionally, the recent incorporation of chain-of-thought reasoning Zhai (2024); Khaliq et al. (2024) has enhanced its ability to support complex problem solving and logical inference. Finally, integrating multimodal RAG into interactive agents and AI assistants such as Gemini Team et al. (2024) enables natural language-driven visual search, document understanding, and multimodal reasoning. The taxonomy of application domains can be seen in Figure 3. The following sections explore domain-specific adaptations of these techniques in greater depth.
Healthcare and Medicine
Multimodal RAG enhances clinical decision-making through integrated analysis of medical imaging, electronic health records, and biomedical literature. Systems like MMED-RAG Xia et al. (2024a) address diagnostic uncertainty in medical visual question answering by aligning radiology images with contextual patient data. RULE Xia et al. (2024b) mitigates hallucinations in automated report generation through dynamic retrieval of clinically similar cases. AsthmaBot Bahaj and Ghogho (2024) introduces a multimodal RAG-based approach for supporting asthma patients across multiple languages, enabling structured, language-specific semantic searches. Predictive frameworks such as Realm Zhu et al. (2024c) demonstrate robust risk assessment by fusing heterogeneous patient data streams, while Hybrid RAG Su et al. (2024a) advances privacy-preserving architectures for federated clinical data integration. FactMM-RAG Sun et al. (2024b) automates radiology report drafting by retrieving biomarker correlations from medical ontologies, exemplifying the paradigm’s capacity to operationalize expert knowledge at scale.
Software Engineering
Code generation systems leverage multimodal RAG to synthesize context-aware solutions from technical documentation and version histories. DocPrompting Zhou et al. (2023) improves semantic coherence in code completion by retrieving API specifications and debugging patterns. Commit message generation models like RACE Shi et al. (2022) contextualize code diffs against historical repository activity, while CEDAR Nashid et al. (2023) optimizes few-shot learning through retrieval-based prompt engineering. REDCODER Parvez et al. (2021) enhances code summarization via semantic search across open-source repositories, preserving syntactic conventions across programming paradigms.
Fashion and E-Commerce
Cross-modal alignment drives advancements in product discovery and design automation. UniFashion Zhao et al. (2024) enables style-aware retrieval by jointly embedding garment images and textual descriptors, while Dang (2024) reduces search friction through multimodal query expansion. LLM4DESIGN Chen et al. (2024c) demonstrates architectural design automation by retrieving compliance constraints and environmental impact assessments, underscoring RAG’s adaptability to creative domains.
Entertainment and Social Computing
Multimedia analytics benefit from RAG’s capacity to correlate heterogeneous signals. SoccerRAG Strand et al. (2024) derives tactical insights by linking match footage with player statistics. MMRA Zhong et al. (2024) predicts content virality through joint modeling of visual aesthetics and linguistic engagement patterns.
Emerging Applications
Autonomous systems adopt multimodal RAG for explainable decision-making, as seen in RAG-Driver’s Yuan et al. (2024) real-time retrieval of traffic scenarios during navigation. ENWAR Nazar et al. (2024) enhances wireless network resilience through multi-sensor fusion, while Riedler and Langer (2024) streamline equipment maintenance by retrieving schematics during fault diagnosis. Geospatial systems such as Img2Loc Zhou et al. (2024e) advance image geolocalization through cross-modal landmark correlation.
6 Open Problems and Future Directions
Despite rapid advancements in multimodal RAG systems, fundamental challenges remain in achieving robust, efficient, and human-like reasoning across modalities.
Generalization, Explainability, and Robustness
Multimodal RAG systems often struggle with domain-specific training and exhibit modality biases, frequently over-relying on text for both retrieval and generation Winterbottom et al. (2020). Explainability remains a key challenge, as these systems typically fail to attribute answers to precise sources. Current methods often cite entire documents or large visual regions as source attribution rather than identifying the exact part of an image, speech, or other modality that led to the answer Ma et al. (2024b); Hu et al. (2023). In addition, the interplay between modalities affects the quality of outcomes produced by these models; for example, answers derived solely from text sources may differ in quality compared to those requiring a combination of text and image inputs Baltrusaitis et al. (2019). They are also vulnerable to adversarial perturbations, such as misleading images influencing textual outputs, and their performance can degrade when relying on low-quality or outdated sources Chen et al. (2022b). While the trustworthiness of unimodal RAGs has been studied Zhou et al. (2024d), enhancing the robustness of multimodal RAGs remains an open challenge and a promising research direction.
Reasoning, Alignment, and Retrieval Enhancement
Multimodal RAGs struggle with compositional reasoning, where information from different modalities must be logically integrated to generate coherent and contextually rich outputs. While cross-modal techniques such as Multimodal-CoT have been proposed Zhang et al. (2023), further innovations are needed to improve the coherence and contextual relevance of multimodal outputs. Enhancing modality alignment and retrieval strategies, particularly for entity-aware retrieval, is essential. Moreover, despite the potential of knowledge graphs to enrich cross-modal reasoning, they remain largely underexplored in multimodal RAGs compared to text-based RAGs Zhang et al. (2024g); Procko and Ochoa (2024). Retrieval biases such as position sensitivity Hu et al. (2024c), redundant retrieval Nan et al. (2024), and biases propagated from training data or retrieved content Zhai (2024), pose significant challenges that require further attention. Another promising direction is developing a unified embedding space for all modalities, enabling direct multimodal search without intermediary conversion models (e.g., ASRs). Despite some progress, mapping multimodal knowledge into a unified space remains an open challenge with significant potential.
Agent-Based and Self-Guided Systems
Recent trends indicate a shift towards agent-based multimodal RAGs that integrate retrieval, reasoning, and generation across diverse domains. Unlike static RAG systems, future multimodal RAGs should integrate interactive feedback and self-guided decision-making to iteratively refine outputs. Furthermore, existing feedback mechanisms often fail to accurately identify whether issues with responses arise during retrieval, generation, or other stages Dong et al. (2024b). The incorporation of reinforcement learning and end-to-end human-aligned feedback into multimodal RAGs remains largely unexplored but holds significant potential for enhancing these systems. These methods could enable multimodal RAGs to assess whether retrieval is necessary, evaluate the relevance of retrieved content, and dynamically determine the most suitable modalities for response generation. Achieving robust support for any-to-any modality is essential for adaptability in open-ended tasks Wu et al. (2024b). Future multimodal RAGs should incorporate data from diverse real-world sources, such as environmental sensors, alongside traditional modalities like text and images, to enhance situational awareness. This progression aligns with the trend toward embodied AI, where models integrate knowledge with physical interaction, enabling applications in robotics, navigation, and physics-informed reasoning. Bridging retrieval-based reasoning with real-world agency brings these systems closer to AGI.
Long-Context Processing, Scalability, and Personalization
High computational costs in video frame sampling and memory bottlenecks in processing multi-page documents with images remain key challenges in long-context processing. Fixed extraction rates often fail to capture relevant video frames, necessitating dynamic adjustment based on content complexity and movement Kandhare and Gisselbrecht (2024). Additionally, retrieval speed-accuracy trade-offs in edge deployments and redundant computations in cross-modal fusion layers highlight the need for more scalable architectures. Personalization mechanisms, such as adapting retrieval to user-specific contexts like medical history, remain underexplored. As personalization mechanisms evolve, ensuring privacy and mitigating the risks of sensitive data leakage in multimodal outputs remain critical challenges. Lastly, the lack of datasets with complex reasoning tasks and multimodal adversarial examples hinders robust evaluation.
7 Conclusion
In this study, we take a deep dive into the realm of multimodal Retrieval-Augmented Generation (RAG) literature. Specifically, we explore and categorize key advancements across different aspects of multimodal RAG systems, including retrieval, multimodal fusion, augmentation, generation, and training strategies. Additionally, we examine the tasks these systems address, their domain-specific applications, and the datasets, benchmarks, and evaluation methods. We also discuss open challenges and limitations in current approaches. We hope it could encourage future research in the domain, particularly in improving cross-modal reasoning and retrieval, developing agent-based interactive systems, and advancing unified multimodal embedding spaces.
8 Limitations
This study offers a comprehensive examination of multimodal RAG systems. Extended discussions, details of datasets and benchmarks, and additional relevant work are available in the Appendices. While we have made our maximum effort; however, some limits may persist. First, due to space constraints, our descriptions of individual methodologies are necessarily concise. Second, although we curate studies from major venues (e.g., ACL, EMNLP, NeurIPS, CVPR, ICLR, ICML, ACM Multimedia) and arXiv, our selection may inadvertently overlook emerging or domain-specific research, with a primary focus on recent advancements. Additionally, this work does not include a comparative performance evaluation of the various models, as task definitions, evaluation metrics, and implementation details vary significantly across studies, and executing these models requires substantial computational resources.
Furthermore, multimodal RAG is a rapidly evolving field with many open questions, such as optimizing fusion strategies for diverse modalities and addressing scalability challenges. As new paradigms emerge, our taxonomy and conclusions will inevitably evolve. To address these gaps, we plan to continuously monitor developments and update this survey and the corresponding GitHub repository to incorporate overlooked contributions and refine our perspectives.
Appendix A Taxonomy
In this section, we present a comprehensive taxonomy of multimodal retrieval-augmented generation (RAG) systems, along with a classification of their application domains. Figures 2 and 3 summarize these taxonomies, respectively.
Figure 2 provides an overview of recent advances in multimodal retrieval-augmented generation (RAG) systems. The taxonomy is organized into several key categories.
  • Retrieval strategies cover efficient search and similarity retrieval methods (including maximum inner product search (MIPS) variants and different multi-modal encoders) and modality-centric techniques that distinguish between text-, vision-, and video-centric as well as document retrieval models. Re-ranking strategies further refine these methods via optimized example selection, relevance scoring, and filtering.
  • Fusion mechanisms are implemented through score fusion and alignment, attention-based techniques, and unified frameworks that project multimodal information into common representations.
  • Augmentation techniques address context enrichment as well as adaptive and iterative retrieval.
  • Generation methods include in-context learning, reasoning, instruction tuning, and source attribution.
  • training strategies are characterized by approaches to alignment and robustness.
Detailed discussions of these categories are provided in the corresponding sections.
Figure 3
presents the taxonomy of application domains for multimodal RAG systems. The identified domains include healthcare and medicine, software engineering, fashion and e-commerce, entertainment and social computing, and emerging applications. This classification offers a concise overview of the diverse applications and serves as a framework for the more detailed analyses that follow.
Appendix B Dataset and Benchmark
Table 1 and Table 2 present a comprehensive overview of datasets and benchmarks commonly employed in multimodal RAG research. The table is organized into five columns:
  • Category: This column categorizes each dataset or benchmark based on its primary domain or modality. The datasets are grouped into eight categories: Image–Text General, Video–Text, Audio–Text, Medical, Fashion, 3D, Knowledge & QA, and Other. The benchmarks are grouped into three categories: Vision-Focused, Cross-Modal and Reasoning-Focused, and Text and Knowledge-Focused. This classification facilitates a clearer understanding of each dataset or benchmark’s role within a multimodal framework.
  • Name: The official name of the dataset or benchmarks is provided along with a citation for reference.
  • Statistics and Description: This column summarizes key details such as dataset size, the nature of the content (e.g., image–text pairs, video captions, QA pairs), and the specific tasks or applications for which the dataset or benchmarks are used. These descriptions are intended to convey the dataset’s scope and its relevance to various multimodal RAG tasks.
  • Modalities: The modalities covered by each dataset or benchmark are indicated (e.g., Image, Text, Video, Audio, or 3D). Notably, several datasets are unimodal; however, within multimodal RAG systems, these are combined with others to represent distinct aspects of a broader multimodal context.
  • Link: A hyperlink is provided to direct readers to the official repository or additional resources for the dataset or benchmark, thereby facilitating further exploration of its properties and applications.
Table 1: Overview of Popular Datasets in Multimodal RAG Research.
Category
Name Statistics and Description Modalities Link
Image-Text General LAION-400M Schuhmann et al. (2021) 200M image–text pairs; used for pre-training multimodal models. Image, Text LAION-400M
Conceptual-Captions (CC) Sharma et al. (2018) 15M image–caption pairs; multilingual English–German image descriptions. Image, Text Conceptual Captions
CIRR Liu et al. (2021) 36,554 triplets from 21,552 images; focuses on natural image relationships. Image, Text CIRR
MS-COCO Lin et al. (2014) 330K images with captions; used for caption–to–image and image–to–caption generation. Image, Text MS-COCO
Flickr30K Young et al. (2014) 31K images annotated with five English captions per image. Image, Text Flickr30K
Multi30K Elliott et al. (2016) 30k German captions from native speakers and human–translated captions. Image, Text Multi30K
NoCaps Agrawal et al. (2019) For zero–shot image captioning evaluation; 15K images. Image, Text NoCaps
Laion-5B Schuhmann et al. (2022) 5B image–text pairs used as external memory for retrieval. Image, Text LAION-5B
COCO-CN Author and Author (2018) 20,341 images for cross-lingual tagging and captioning with Chinese sentences. Image, Text COCO-CN
CIRCO Baldrati et al. (2023) 1,020 queries with an average of 4.53 ground truths per query; for composed image retrieval. Image, Text CIRCO
Video-Text BDD-X Xu et al. (2018) 77 hours of driving videos with expert textual explanations; for explainable driving behavior. Video, Text BDD-X
YouCook2 Zhou et al. (2018) 2,000 cooking videos with aligned descriptions; focused on video–text tasks. Video, Text YouCook2
ActivityNet Caba Heilbron et al. (2015) 20,000 videos with multiple captions; used for video understanding and captioning. Video, Text ActivityNet
SoccerNet Giancola et al. (2018) Videos and metadata for 550 soccer games; includes transcribed commentary and key event annotations. Video, Text SoccerNet
MSR-VTT Xu et al. (2016) 10,000 videos with 20 captions each; a large video description dataset. Video, Text MSR-VTT
MSVD Chen and Dolan (2011) 1,970 videos with approximately 40 captions per video. Video, Text MSVD
LSMDC Rohrbach et al. (2015) 118,081 video–text pairs from 202 movies; a movie description dataset. Video, Text LSMDC
DiDemo Anne Hendricks et al. (2017) 10,000 videos with four concatenated captions per video; with temporal localization of events. Video, Text DiDemo
Breakfast Kuehne et al. (2014) 1,712 videos of breakfast preparation; one of the largest fully annotated video datasets. Video, Text Breakfast
COIN Tang et al. (2019) 11,827 instructional YouTube videos across 180 tasks; for comprehensive instructional video analysis. Video, Text COIN
MSRVTT-QA Xu et al. (2017) Video question answering benchmark. Video, Text MSRVTT-QA
MSVD-QA Xu et al. (2017) 1,970 video clips with approximately 50.5K QA pairs; video QA dataset. Video, Text MSVD-QA
ActivityNet-QA Yu et al. (2019) 58,000 human–annotated QA pairs on 5,800 videos; benchmark for video QA models. Video, Text ActivityNet-QA
EpicKitchens-100 Dima (2020) 700 videos (100 hours of cooking activities) for online action prediction; egocentric vision dataset. Video, Text EPIC-KITCHENS-100
Ego4D Grauman et al. (2022) 4.3M video–text pairs for egocentric videos; massive–scale egocentric video dataset. Video, Text Ego4D
HowTo100M Miech et al. (2019) 136M video clips with captions from 1.2M YouTube videos; for learning text–video embeddings. Video, Text HowTo100M
CharadesEgo Sigurdsson et al. (2018) 68,536 activity instances from ego–exo videos; used for evaluation. Video, Text Charades-Ego
ActivityNet Captions Krishna et al. (2017) 20K videos with 3.7 temporally localized sentences per video; dense–captioning events in videos. Video, Text ActivityNet Captions
VATEX Wang et al. (2019) 34,991 videos, each with multiple captions; a multilingual video–and–language dataset. Video, Text VATEX
Charades Sigurdsson et al. (2016) 9,848 video clips with textual descriptions; a multimodal research dataset. Video, Text Charades
WebVid Bain et al. (2021) 10M video–text pairs (refined to WebVid-Refined-1M). Video, Text WebVid
Youku-mPLUG Xu et al. (2023) Chinese dataset with 10M video–text pairs (refined to Youku-Refined-1M). Video, Text Youku-mPLUG
Audio-Text LibriSpeech Panayotov et al. (2015) 1,000 hours of read English speech with corresponding text; ASR corpus based on audiobooks. Audio, Text LibriSpeech
AudioCap Kim et al. (2019) 46K audio clips paired with human-written text captions. Audio, Text AudioCaps
AudioSet Gemmeke et al. (2017) 2,084,320 human–labeled 10–second sound clips from YouTube; 632 audio event classes. Audio, Text AudioSet
Medical MIMIC-CXR Johnson et al. (2019) 125,417 training pairs of chest X–rays and reports. Image, Text MIMIC-CXR
CheXpert Irvin et al. (2019) 224,316 chest radiographs of 65,240 patients; focused on medical analysis. Image, Text CheXpert
MIMIC-III Johnson et al. (2016) Health-related data from over 40K patients (text data). Text MIMIC-III
IU-Xray Pavlopoulos et al. (2019) 7,470 pairs of chest X–rays and corresponding diagnostic reports. Image, Text IU X-ray
PubLayNet Zhong et al. (2019) 100,000 training samples and 2,160 test samples built from PubLayNet (tailored for the medical domain). Image, Text PubLayNet
Fashion Fashion-IQ Wu et al. (2019) 77,684 images across three categories; evaluated with Recall@10 and Recall@50. Image, Text Fashion IQ
FashionGen Hadi Kiapour et al. (2018) 260.5K image–text pairs of fashion images and item descriptions. Image, Text Fashion-Gen
VITON-HD Choi et al. (2021) 83K images for virtual try–on; high–resolution clothing items. Image, Text VITON-HD
Fashionpedia Author and Author (2023a) 48,000 fashion images annotated with segmentation masks and fine-grained attributes. Image, Text Fashionpedia
DeepFashion Liu et al. (2016) Approximately 800K diverse fashion images for pseudo triplet generation. Image, Text DeepFashion
3D ShapeNet Chang et al. (2015) 7,500 text–3D data pairs; repository for 3D CAD models. Text, 3D ShapeNet
Knowledge & QA VQA Antol et al. (2015) 400K QA pairs with images for visual question answering. Image, Text VQA
PAQ Lewis et al. (2021) 65M text–based QA pairs; a large–scale dataset. Text PAQ
ELI5 Fan et al. (2019) 270K complex and diverse questions augmented with web pages and images. Text ELI5
ViQuAE Biten et al. (2022) 11.8M passages from Wikipedia covering 2,397 unique entities; knowledge–intensive QA. Text ViQuAE
OK-VQA Marino et al. (2019) 14K questions requiring external knowledge for VQA. Image, Text OK-VQA
WebQA Li et al. (2022b) 46K queries that require reasoning across text and images. Text, Image WebQA
Infoseek Li et al. (2021) Fine-grained visual knowledge retrieval using a Wikipedia–based knowledge base ( 6M passages). Image, Text Infoseek
ClueWeb22 Callan et al. (2022) 10 billion web pages organized into three subsets; a large–scale web corpus. Text ClueWeb22
MOCHEG Yao et al. (2023) 15,601 claims annotated with truthfulness labels and accompanied by textual and image evidence. Text, Image MOCHEG
VQA v2 Goyal et al. (2017b) 1.1M questions (augmented with VG–QA questions) for fine-tuning VQA models. Image, Text VQA v2
ScienceQA Lu et al. (2022) High–school science curriculum dataset with 12K training samples and over 21K questions across various subjects. Text ScienceQA
A-OKVQA Schwenk et al. (2022) Benchmark for visual question answering using world knowledge; around 25K questions. Image, Text A-OKVQA
MMBench Liu et al. (2025) Bilingual benchmark for assessing the multimodal capabilities of VLMs; over 3,000 multiple–choice questions covering 20 dimensions. Text MMBench
XL-HeadTags Shohan et al. (2024) 415K news headline-article pairs consist of 20 languages across six diverse language families. Text XL-HeadTags
SEED-Bench Li et al. (2023a) 19K multiple–choice questions with accurate human annotations across 12 evaluation dimensions. Text SEED-Bench
Other ImageNet Deng et al. (2009) 14,197,122 images for perspective understanding; a hierarchical image database. Image ImageNet
Oxford Flowers102 Nilsback and Zisserman (2008) 102 flower categories with five examples per category; image classification dataset. Image Oxford Flowers102
Stanford Cars Krause et al. (2013) Images of different car models (five examples per model); for fine-grained categorization. Image Stanford Cars
GeoDE Author and Author (2023b) 61,940 images from 40 classes across 6 world regions; emphasizes geographic diversity in object recognition. Image GeoDE
Table 2: Overview of Popular Benchmarks in Multimodal RAG Research.
Category Name Statistics and Description Modalities Link
Vision-Focused MRAG-Bench Hu et al. (2024c)
Evaluates visual information retrieval and integration;
tests robustness to irrelevant visual information and visual reasoning.
Images MRAG-Bench
VQAv2 Goyal et al. (2017a)
Contains 265,016 images with 1,105,904 questions; assesses the ability to answer questions about images,
integrating image and question text, visual reasoning, and common-sense reasoning.
Images + Text VQAv2
OK-VQA Marino et al. (2019)
Comprises 14,031 questions; tests retrieval and integration of external information,
combining image, question, and external knowledge, focusing on reasoning with external knowledge.
Images + Text OK-VQA
Cross-Modal and Reasoning-Focused M2RAG Ma et al. (2024c)
Establishes a comprehensive benchmark for multimodal retrieval-augmented generation;
integrates fine-grained text-modal and multimodal metrics to evaluate foundation models on dynamic retrieval,
multi-hop reasoning, and multimodal integration.
Images + Text M2RAG
Dyn-VQA Li et al. (2024b)
Emphasizes dynamic retrieval with rapidly changing information;
requires integration of dynamic knowledge retrieval and multi-hop reasoning, testing robustness to changing information.
Images + Text Dyn-VQA
MMBench Liu et al. (2025)
Covers various tasks, including visual question answering, image captioning, and cross-modal retrieval;
focuses on integrating information across multiple modalities (vision, text, audio) for general multimodal understanding.
Images + Text + Audio MMBench
ScienceQA Saikh et al. (2022)
Contains 21,208 questions; tests scientific reasoning abilities using multimodal contexts (text, diagrams, images);
requires deep understanding and integration of information, scientific reasoning, and diagram understanding.
Images + Diagrams + Text ScienceQA
SK-VQA Su et al. (2024c)
Offers over 2 million question-answer pairs;
focuses on synthetic knowledge generation, multimodal reasoning, and external knowledge integration;
used for training context-augmented multimodal large language models.
Images + Text SK-VQA
Text and Knowledge-Focused TriviaQA Joshi et al. (2017)
Provides 650K question-answer pairs; a reading comprehension dataset where questions are based on evidence documents;
can be adapted to multimodal RAG by providing images or other visual context related to the trivia;
focuses on knowledge retrieval and factual accuracy.
Text TriviaQA
Natural Questions Kwiatkowski et al. (2019)
Contains 307,373 training examples; questions are naturally occurring Google search queries;
adapting this for multimodal RAG involves finding or creating relevant visual contexts for the queries;
emphasizes knowledge retrieval and handling noisy or ambiguous queries.
Text Natural Questions
SMMQG Wu et al. (2024a)
Includes 1,024 question-answer pairs; focuses on synthetic data generation,
multimodal retrieval-augmented generation, and control over question style and modality.
Images + Text SMMQG
Appendix C Metrics
Accuracy: Accuracy is typically defined as the ratio of correctly predicted instances to the total instances. In retrieval-based tasks, Top-K Accuracy is defined as:
[…]
FID (Fréchet inception distance): FID is a metric used to assess the quality of images created by a generative model. The formula for FID is:
FID $=\|\mu_{r}-\mu_{g}\|^{2}+tr(\Sigma_{r}+\Sigma_{g}-2\sqrt{\Sigma_{r}\Sigma_{g}})$ (6)
where $\mu_{r}$ and $\Sigma_{r}$ are the mean and covariance of real images’ feature representations, respectively. $\mu_{g}$ and $\Sigma_{g}$ are the mean and covariance of generated images’ feature representations, respectively. To extract features, InceptionV3 is typically used.
ROUGE-N (N-gram Recall): The ROUGE metric is commonly used to evaluate text summarization and generation. ROUGE-N measures the overlap of N-grams between the generated and reference text. The formula for ROUGE-N is:
ROUGE-N $=\frac{\sum_{\text{gram}_{N}\in\text{Ref}}\text{Count}_{\text{match}}(\text{gram}_{N})}{\sum_{\text{gram}_{N}\in\text{Ref}}\text{Count}(\text{gram}_{N})}$ (7)
ROUGE-L measures the longest common subsequence (LCS) between generated and reference text. The formula for ROUGE-L is:
ROUGE-L $=\frac{LCS(X,Y)}{|Y|}$ (8)
BLEU: BLEU is another metric used for assessing text generation. The formula for calculating BLEU is:
$\text{BLEU}(p_{n},\text{BP})=\text{BP}\cdot\exp\left(\sum_{n=1}^{N}w_{n}\log p_{n}\right)$ (9)
Here, $p_{n}$ represents the precision of n-grams, $w_{n}$ denotes the weight assigned to the n-gram precision, and the Brevity Penalty (BP) is defined as:
$\text{BP}=\begin{cases}1&\text{length}>\text{$rl$}\\ \exp\left(1-\frac{\text{$rl$}}{\text{$cl$}}\right)&\text{length}\leq\text{$rl$}\end{cases}$ (10)
Here, $rl$ represents the reference length and $cl$ represents the candidate length.
BERTScore: BERTScore is a metric for evaluating the quality of text generation, based on the similarity between the contextual embeddings of words in the candidate and reference texts. The formula for calculating BERTScore is:
$\text{BERTScore}(c,r)=\frac{1}{|c|}\sum_{i=1}^{|c|}\max_{j=1}^{|r|}\cos(\mathbf{e}_{i},\mathbf{e}_{j})$ (11)
$c$ is the candidate sentence, and $r$ is the reference sentence. $e_{i}$ and $e_{j}$ are the embeddings (e.g., from BERT) for words $c_{i}$ and $r_{j}$ in the candidate and reference sentences, respectively.
CLIPScore: CLIPScore is a metric that evaluates the similarity between the text and an image by using the CLIP model. The formula for calculating CLIPScore is:
$\text{CLIPScore}=\frac{\cos(\mathbf{t},\mathbf{i})}{\|\mathbf{t}\|\cdot\|\mathbf{i}\|}$ (12)
where t and i are text and image embedding respectively.
Mean Reciprocal Rank (MRR): MRR is a metric used to evaluate the performance of systems that return a list of results, such as search engines or recommendation systems. MRR measures the rank position of the first relevant result in the returned list. The formula for calculating MRR is:
$\text{MRR}=\frac{1}{Q}\sum_{q=1}^{Q}\frac{1}{\text{rank}_{q}}$ (13)
where $Q$ is the total number of queries. $rank_{q}$ is the rank of the first relevant result for query $q$.
Appendix D Loss Function
InfoNCE (Information Noise Contrastive Estimation): The InfoNCE loss is commonly used in self-supervised learning, especially in contrastive learning methods. The formula for InfoNCE loss is:
$\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp(\text{sim}(z_{i},z_{j})/\tau)}{\sum_{k=1}^{K}\exp(\text{sim}(z_{i},z_{k})/\tau)}$ (14)
where $z_{i}$ and $z_{j}$ are the embeddings of a positive pair and $\tau$ is a temperature parameter.
GAN (Generative Adversarial Network): The GAN loss consists of two parts: the discriminator loss and the generator loss. The discriminator loss formula is:
$\scriptstyle\mathcal{L}_{D}=-\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D(x)]-\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))]$ (15)
where $x$ is a real sample from the data distribution. $G(z)$ is the generated sample from the generator, where $z$ is a noise vector. $D(x)$ is the probability that $x$ is real.
The Generator loss formula is:
$\mathcal{L}_{G}=-\mathbb{E}_{z\sim p_{z}(z)}[\log D(G(z))]$ (16)
Triplet Loss: Triplet Loss is used in metric learning to ensure that similar data points are closer together while dissimilar ones are farther apart in the embedding space. The Triplet loss formula is:
$\scriptstyle\mathcal{L}=\sum_{i=1}^{N}\max(0,\|f(x_{a}^{i})-f(x_{p}^{i})\|^{2}-\|f(x_{a}^{i})-f(x_{n}^{i})\|^{2}+\alpha)$ (17)
where $x_{a}^{i}$ is the anchor sample. $x_{p}^{i}$ and $x_{n}^{i}$ are the positive and negative samples respectively. $f(x)$ is the neural network.